The length of a rectangle is 11 ft. longer than twice the width. If the perimeter is 94 ft, find the length and width of the rectangle.
Answer by Cromlix(4381)

Hi there
Make the width = 'x'
So, the length = 11 + 2x
Perimeter = 2 x Length + 2 x Width
94 = 2(11 + 2x) + 2x
94 = 22 + 4x + 2x
Collect like terms
6x = 94 - 22
6x = 72
x = 12
Width = 12 ft
Length = 35 ft
Hope this helps:-)